# Leveraging Moving Sound Source Trajectories for Universal Sound Separation

### v1

@@ III-A Initial tracking
After the envelope is estimated, it is concatenated with the audio mixture and input to the tracking network. The network used is our previously proposed DCSAnet [30], whose architecture is illustrated in Fig 2. A 1-D convolutional layer with learnable parameters is used to encode input signals followed by dilated convolutional layers [31]. After that a multi-head self-attention layer is utilized to capture the global temporal information. Finally, a 1-D deconvolution decoder is used to convert the features of network to temporal domain. The output of the tracking network is denoted as $traj^{est}\in R^{3\times T}$, where $T$ is the number of temporal samples of the trajectory which is equal to the length of input audio signal. The 3D vector at each temporal sample is supposed to be the XYZ coordinates of a unit vector pointing the the sound source.
@@
To mitigate the impact when the sound energy is low and to ensure more accurate tracking results when the sound source is active, we multiply the trajectory and the envelope of source energy and formulate a trajectory of intensity vectors, which we call it intensity trajectory, denoted as $I\_traj\in R^{3\times T}$. At each temporal sample $t$, the 3D vector simultaneously represents the energy and direction of the sound source. The length of the vector indicates the energy of the sound source, while the direction of the vector points to the location of the sound source. The loss function for training the tracking network consists of two parts: the intensity trajectory loss and the differential loss:
@@
$\begin{split}Loss_{track}&=\alpha Loss_{traj}+\beta Loss_{diff},\end{split}$ (3)
@@
$\begin{split}Loss_{traj}&=MSELoss(E^{est}\cdot traj^{est},E^{est}\cdot traj^{trg})\\ &=\frac{1}{T}\sum_{t=0}^{T}(I\_traj^{est}[t]-I\_traj^{trg}[t])^{2},\end{split}$ (4)
where $traj^{trg}\in R^{3\times T}$ is the XYZ coordinates of the unit vector pointing to the sound source at each temporal sample. The definition of differential loss is the same as [30].
@@ III-B Mutual facilitation
The mutual facilitation module includes two networks: the target sound extraction network and the precise tracking network. The network used for target sound extraction in this study is also SpatialNet with Mamba [23]. In this paper, we modify SpatialNet to serve as the target sound extraction network. The mixed signal is transformed to T-F domain by Short Time Fourier Transform (STFT), which is denoted as $\mathbf{Y}\in\mathbb{R}^{2M\times N\times F}$, where $M=4$ is the number of orders of the FOA and $N$ is the number of frames and $F$ is the number of frequency points. The intensity trajectory of the target sound source is also converted into frames by averaging the intensity trajectory in one frame, which can be denoted as $\mathbf{I\_Traj}\in\mathbb{R}^{3\times N\times 1}$. The T-F bins of the mixture and frames of the intensity trajectory each undergo a different 1D convolution encoder to obtain features, which can be denoted as $\mathbf{Y_{f}}\in\mathbb{R}^{H\times N\times F}$ and $\mathbf{I\_Traj_{f}}\in\mathbb{R}^{H\times N\times 1}$, where $H$ is the size of hidden state. The signal feature and the trajectory feature are then multiplied together and fed into subsequent cross-band and narrow-band modules. Unlike the original SpatialNet, where the network’s output was the single-channel signals for all sound sources, the modified SpatialNet now outputs the FOA multi-channel signal of the target sound source. The architecture of the modified SpatialNet is shown in Fig 3.
@@
During the training of the target sound extraction network, we do not use the output of the initial tracking module as the input. Instead, the target intensity trajectory $I\_traj^{trg}$ is utilized as input. This aims to establish the performance upper bound of the target sound extraction network, ensuring that its separation performance improves with the precision of the intensity trajectory. Therefore, the training of the target sound extraction network is independent of the initial tracking module. The loss function for training the target sound extraction network is the negative signal-to-noise-ratio (SNR), which is the same as SpatialNet.
@@
After the multi-channel target sound is extracted, we calculate its energy envelope, concatenate the extracted multi-channel target signal and the calculated envelope along with the signal mixture and feed them into the precise tracking. The precise tracking module is based on the separated signal thus can generate accurate intensity trajectories. The network structure and loss function for precise tracking are the same as those for initial tracking, except that the number of the input channels of the 1D convolutional layer is set to $9$ instead of $5$. During training, the output of the target sound extraction network when inputting the correct intensity trajectory label is utilized as input. The purpose of this approach is also to obtain the performance upper bound of precise tracking, ensuring that its tracking performance improves with the enhancement of separation performance.
@@
During testing, we first estimate the energy envelopes of all the sound sources based on the signal mixture $Y\in R^{M\times T}$, select one of them, which is denoted as $E^{est}\in R^{1\times T}$, and perform initial tracking based on the $E^{est}$ and signal mixture. The initial intensity trajectory denoted as $I\_traj_{init}$ is the starting condition of mutual facilitation. We feed it into the target sound extraction network and the separated multi-channel signal of the target source is extracted, which is denoted as $X_{sep1}\in R^{M\times T}$. The $X_{sep1}$ assists the precise tracking with $E^{est}$ and $Y$ and a more precise intensity trajectory $I\_traj_{prec1}$ is generated. Based on this precise intensity trajectory, we can perform the next round of mutual facilitation, thus a more accurate multi-channel signal $X_{sep2}$ and a more precise intensity trajectory $I\_traj_{prec2}$ can be generated. It is worth mentioning that, we don’t change the estimated energy envelope $E^{est}$ during this procedure, as the first extracted multi-channel target sound may not be accurate enough, and the energy envelope calculated based on it can do harm to the following precise tracking. Besides, we also find that the initial estimated energy envelope is accurate enough.
@@ III-C Neural beamforming
The neural beamformer is trained on the more accurate multi-channel signal $X_{sep2}$, the more precise intensity trajectory $I\_traj_{prec2}$ and $Y$ to utilize the multi-channel information and generate a cleaner single-channel signal. In this paper, the zeroth-order signal of the FOA signal is selected as the signal of the reference channel. The network of the neural beamformer is also based on SpatialNet. Its structure is shown in Fig 4. The mixed signals and the multi-channel separated signals are transformed to T-F domain through STFT, and then they are concatenated along the channel dimension, which is denoted as $\mathbf{Y_{cat}}\in\mathbb{R}^{4M\times T\times F}$. The rest of the network structure of this neural beamformer is the same as that for target sound extraction, except that the number of input channels of the 1D convolutional layer for audio input is set to $4M$, and the number of output channels of the 1D deconvolutional layer is reduced to $1$. The neural beamformer still multiplies the framed intensity trajectory and the time-frequency domain signal after the encoders. The loss function of training the neural beamformer follows the one of target sound extraction network.
@@ IV-A Experimental setup
(1) Dataset generation: Our approach is validated using a simulated dataset sourced from FSD18k [32]. The dataset comprises 45 hours of signals sampled at 16kHz for training and 5 hours for testing. Room dimensions are randomly selected from a range of $3\times 3\times 3m^{3}$ to $10\times 8\times 6m^{3}$, with reverberation times $T60$ randomly assigned between $0.2$s and $0.8$s. The microphone array, configured as an Eigenmike microphone array with 32 channels, is positioned within $10\%$ of the room’s size from the center in each dimension. The captured signals are then converted into FOA format.
@@
The method for generating trajectories of the sound sources is same as that outlined in [33]. Initially, two points are randomly selected within the room: a starting point $\mathbf{p_{0}}=(p_{x0},p_{y0},p_{z0})$ and an endpoint $\mathbf{p_{T}}=(p_{xT},p_{yT},p_{zT})$. Connecting these two points establishes the fundamental trajectory line. Subsequently, sinusoidal trajectories with random frequencies $\mathbf{\omega}=(\omega_{x},\omega_{y},\omega_{z})$ and amplitudes $\mathbf{A}=(A_{x},A_{y},A_{z})$ are added along each axis. The number of oscillation periods is constrained to not exceed $2$. The position of the sound source at temporal sample $t$ can be calculated as:
$\begin{split}\mathbf{p_{t}}=\mathbf{p_{0}}+\frac{t}{T}(\mathbf{p_{T}}-\mathbf{p_{0}})+\mathbf{A}\cdot sin(\mathbf{\omega}t),\end{split}$ (5)
@@
where $T$ is the length of the trajectory. In this experiment, only the DOA of the sound sources in utilized for target sound extraction. We set the number of sound sources to $2$. Further research on varying the number of sound sources will be conducted in future work. The RIRs is generated based on the Image Source Method [34] with the gpuRIR Python library [35].
@@
(2) Network configuration: The network architecture utilized is SpatialNet [21], which consists of $8$ blocks. The hidden unit size in each block is $96$. The kernel size of the encoders of signals and trajectories is $5$. The kernel size of time-dimension group convolution and frequency-dimension group convolution are $5$ and $3$, respectively. The group numbers are $8$ for both. We employ the SpatialNet adapted based on the Mamba model [23], which replaces the multi-head attention and T-ConvFFN in the narrowband block with Mamba [22]. In Mamba, the dmodel parameter is set to $32$, state is set to $16$, dconv is set to $4$, and expand is set to $2$. The window length and hop size for maximum filter for calculating the energy envelope is 256 samples (16ms). The STFT window length is 256 samples (16ms), the hop size is 128 samples (8ms), and the number of frequency bins is 129. Each network in this paper is trained with Adam optimizer [36] with learning rate of $45e-$ 4.
@@
(3) Baselines: we compare our method with Fasnet-TAC [10], which is a neural beamformer, and Conv-Tasnet [4]. The multi-channel end to end networks of SpatialNet with the same architecture of the network in this paper is also utilized as baseline.
@@
Method SI-SNR SNR SDR
FasNet-TAC 12.19 12.72 13.78
Conv-Tasnet 11.83 12.50 13.02
SpatialNet 13.66 14.23 14.94
MFTS 15.21 16.07 16.55
@@ IV-B1 Sound separation
The metrics utilized for evaluating the sound separation performance is SNR, signal to distortion ratio (SDR) and scale-invariant-signal-to-noise-ratio (SI-SNR). Table I shows each method’s average performance for source separation on the test dataset. The method proposed in this paper achieves the best performance. Compared to the baseline methods, our approach explicitly tracks the sound sources and utilizes the tracking results to aid separation. The improved separation results further enhance the accuracy of tracking, enabling the use of more precise tracking results to facilitate better separation.
Figure 5: Separation performance under different T60s.
@@
To demonstrate the robustness of our proposed method, we compute the separation performance of different methods at different T60 values. Fig 5 shows the results. It can be seen that our method maintains good performance under different reverberation times, and outperforms the baseline methods in all cases especially when the reverberation time is low.
Figure 6: The EWRMSAE under different T60s.
IV-B2 Ablation study
To demonstrate the effectiveness of each part of the proposed method, we perform the ablation study.
We first check the effectiveness of the envelope estimation network, it shows that the NMSE is below -19dB under all the reverberation times. To illustrate the efficiency of sound source tracking based on energy envelope estimation compared to directly tracking the sources from the audio mixture, we compare the performance results of both methods. We call the approach tracking the source based on the energy envelope ENtrack, and the approach tracking the sources from audio mixture Mixtrack. Both of these two approaches utilize the same network, except that Mixtrack only takes the audio mixture as input, and outputs all the sources’ trajectories without utilizing the energy envelope. The loss function for training Mixtrack consists of the Mean Square Error (MSE) between the estimated trajectories and the differential loss. Utterance-level pit (uPIT) [37] is used during training Mixtrack, which means that we perform the permutation on the entire trajectory and find the order of permutation with minimum value of the loss. The performance of sound source tracking depends on the energy of the sound source; when the energy is low, the localization accuracy tends to deteriorate. On the other hand, the task of sound source separation focuses more on separating the source when it is active. Therefore, good localization results should aim for greater accuracy when the sound source is active, i.e., when the energy of the source is high. Based on this, we propose an Energy-Weighted Root Mean Square Angular Error (EWRMSAE) as an enhancement of the traditional RMSAE. The EWRMSAE is defined as follows:
$\begin{split}&EWRMSAE=\\ &\sqrt{\frac{\sum_{t=0}^{T}(E^{trg}[t])^{2}\cdot(AE(\mathbf{traj}^{est}[t],\mathbf{traj}^{trg}[t]))^{2}}{\sum_{t=0}^{T}(E^{trg}[t])^{2}}},\end{split}$ (6)
where $AE(\cdot)$ calculates the absolute value of the angular error between the estimated and target trajectory at each temporal sample $t$. EWRMSAE assigns more weights to angular errors when sources have higher energy, while the weights of the angular errors when sources have lower energy is less. Fig 6 shows the EWRMSAE for both ENtrack and Mixtrack under different T60s. It can be observed that tracking based on energy envelope results in a lower EWRMSAE across different T60 values compared to tracking that does not use energy envelope information.
@@ IV-B2 Ablation study
Figure 7: The angular error (AE) at individual audio samples under different amplitudes of the energy envelope.

### v2

@@ III-A2 Initial tracking
After the envelope signal of the $c$-th source $\mathbf{E}^{est}_{c}$ is estimated, it is concatenated with the audio mixture and input to the tracking network to track the $c$-th source. The network used is our previously proposed DCSAnet [39], whose architecture is illustrated in Fig. 4. A 1-D convolutional layer with learnable parameters is used to encode input signals followed by dilated convolutional layers [40]. After that a multi-head self-attention layer is utilized to capture the global temporal information. Finally, a 1-D deconvolution decoder is used to convert the features of network to temporal domain. The output of the tracking network is denoted as $\mathbf{traj}^{est}_{c}\in\mathbb{R}^{N\times 3}$, where $N$ is the number of temporal samples of the trajectory which is equal to the length of input audio signal. The 3-D vector at each temporal sample is supposed to be the XYZ coordinates of a unit vector pointing the the sound source.
@@
To mitigate the impact when the sound energy is low and to ensure more accurate tracking results when the sound source is active, we multiply the trajectory and the envelope signal and formulate a trajectory of intensity vectors, which we call it intensity trajectory, denoted as $\mathbf{I\_traj}_{c}\in\mathbb{R}^{N\times 3}$. At each temporal sample $n$, the 3-D vector simultaneously represents the energy and direction of the sound source. The length of the vector indicates the energy of the sound source, while the direction of the vector points to the location of the sound source. This concept is similar to ACCDOA in Sound event localization and detection (SELD) tasks [41],[42], but there are some differences. In ACCDOA, the vector’s length indicates the estimated likelihood of the source’s presence. In contrast, in the proposed intensity trajectory, the vector’s length reflects the source’s energy. The loss function for training the tracking network consists of two parts: the intensity trajectory loss and the differential loss:
@@
$\begin{split}\mathcal{L}_{track}&=\alpha\cdot\mathcal{L}_{traj}+\beta\cdot\mathcal{L}_{diff}.\end{split}$ (3)
@@
=$\begin{split}\mathcal{L}&{}_{traj}=\text{MSELoss}(\mathbf{I\_traj}^{est}_{c},\mathbf{I\_traj}^{trg}_{c})\\ &=\frac{1}{N}\sum_{n=0}^{N}(E^{est}_{c}(n)\cdot traj^{est}_{c}(n)-E^{est}_{c}(n)\cdot traj^{trg}_{c}(n))^{2},\end{split}$ (4)
where $\mathbf{traj}^{trg}_{c}\in\mathbb{R}^{N\times 3}$ is the XYZ coordinates of the unit vector pointing to the sound source at each temporal sample. The definition of differential loss is the same as [39], which minimizes the error between the estimated and the target trajectory’s average speed at different time scales. The differential of trajectory at time scale $d$ and sample $n$ is defined as:
$\begin{split}\Delta{traj}_{d,c}(n)=traj_{c}(n)-traj_{c}(n-d).\end{split}$ (5)
The differential loss can be calculated as:
=$\begin{split}\mathcal{L}&{}_{diff,d}=\text{MSELoss}(\Delta\mathbf{traj}_{d,c}^{est},\Delta\mathbf{traj}_{d,c}^{trg})\\ &=\frac{1}{N-d+1}\sum_{n=0}^{N-d+1}(\Delta traj^{est}_{d,c}(n)-\Delta traj^{trg}_{d,c}(n))^{2},\end{split}$ (6)
$\begin{split}\mathcal{L}_{diff}=\frac{1}{D}\sum_{i=0}^{D}\mathcal{L}_{diff,2^{i}},\end{split}$ (7)
where $\Delta\mathbf{traj}_{d,c}^{est}$ and $\Delta\mathbf{traj}_{d,c}^{est}$ respectively denote the differential of the estimated and target trajectories at time scale $d$. We choose $D=10$, which is the same as [39].
@@ III-B Mutual facilitation
The mutual facilitation module includes two networks: the target sound extraction network and the precise tracking network. We will first introduce the training of these two networks and then explain the mutual facilitation process during the test phase.
III-B1 Target sound extraction
@@
The network used for target sound extraction in this study is also based on SpatialNet with Mamba [25]. In this paper, we modify SpatialNet to serve as the target sound extraction network. The T-F domain representation $\mathbf{Y}\in\mathbb{R}^{F\times T\times 2M}$ of the audio mixture first undergoes a 1-D convolutional layer to obtain signal features $\mathbf{Y_{f}}\in\mathbb{R}^{F\times T\times H}$, where $H$ is the size of hidden state. The intensity trajectory of the target sound source is also converted into frames by averaging the intensity trajectory in one frame, which can be denoted as $\mathbf{I\_Traj}_{c}\in\mathbb{R}^{1\times T\times 3}$. Then the framed intensity trajectory undergoes a different 1-D convolution encoder to obtain trajectory features, which can be denoted as $\mathbf{I\_Traj\_f}_{c}\in\mathbb{R}^{1\times T\times H}$. The signal feature and the trajectory feature are then multiplied together and fed into subsequent cross-band and narrow-band modules. Unlike the original SpatialNet, where the network’s output was the single-channel signals for all sound sources, the modified SpatialNet now outputs the FOA multi-channel signal of the target sound source. The architecture of the modified SpatialNet is shown in Fig. 5.
@@
In order to establish the performance upper bound of the target sound extraction network, ensuring that its separation performance improves with the precision of the intensity trajectory, unlike the testing phase, during the training of the target sound extraction network, we do not use the output of the initial tracking module as the input. Instead, the groundtruth intensity trajectory $\mathbf{I\_traj}^{trg}_{c}$ is utilized as input. Therefore, the training of the target sound extraction network is independent of the initial tracking module. The loss function for training the target sound extraction network is the negative signal-to-noise-ratio (SNR), which is the same as SpatialNet [23].
III-B2 Precise tracking
@@
After training completion of the target sound extraction network, we utilize its multi-channel output signal to compute the single-channel envelope, concatenate the multi-channel signal and the calculated envelope signal along with the signal mixture and feed them into the precise tracking network. The precise tracking module is based on the cleaner separated signal thus can generate precise intensity trajectories. The network structure and loss function for precise tracking are the same as those for initial tracking, except that the number of the input channels of the 1-D convolutional layer is set to $12M+$ 1 ($M$ for extracted multi-channel target signal, $M$ for the signal mixture and $1$ for the estimated envelope signal of the target sound).
Similar to the training procedure of the target sound extraction network, during training the precise tracking network, the output of the target sound extraction network when inputting the groundtruth intensity trajectory is utilized as input, which constitutes the optimal input attainable by the precise tracking network. The purpose of this approach is also to obtain the performance upper bound of precise tracking, ensuring that its tracking performance improves with the enhancement of separation performance.
III-B3 Mutual facilitation
@@
The mutual facilitation process in utilized in the testing phase. During testing, we first estimate the envelope signals of all the sound sources based on the signal mixture, select one of them, which is denoted as $\mathbf{E}^{est}_{c}\in\mathbb{R}^{N\times 1}$, and perform initial tracking based on the $\mathbf{E}^{est}_{c}$ and signal mixture. The initial coarse intensity trajectory denoted as $\mathbf{I\_traj}_{init,c}$ is the starting clue of mutual facilitation. We feed it into the target sound extraction network and the separated multi-channel signal of the target source is extracted, which is denoted as $\mathbf{X}_{sep1,c}\in\mathbb{R}^{N\times M}$. The $\mathbf{X}_{sep1,c}$ assists the precise tracking with $\mathbf{E}^{est}_{c}$ as well as the audio mixture and then a precise intensity trajectory $\mathbf{I\_traj}_{prec1,c}$ is generated. Based on this precise intensity trajectory, we can perform the next round of mutual facilitation, thus the refined multi-channel signal $\mathbf{X}_{sep2,c}$ and intensity trajectory $\mathbf{I\_traj}_{prec2,c}$ can be generated. It is worth mentioning that, we do not change the estimated envelope signal $\mathbf{E}^{est}_{c}$ during this procedure, as the first extracted multi-channel target sound $\mathbf{X}_{sep1,c}$ may not be accurate enough, and the envelope signal calculated based on it can do harm to the following precise tracking. Besides, we also find that the initially estimated envelope signal proves sufficiently accurate, and updating it in later stages yields no performance benefits while increasing redundancy. This mutual facilitation process is performed for each sound source based on the estimated envelopes.
@@ III-C Neural beamforming
The neural beamformer is trained on the refined multi-channel signal $\mathbf{X}_{sep2,c}$ and intensity trajectory $\mathbf{I\_traj}_{prec2,c}$, as well as the signal mixture to utilize the multi-channel information and generate a cleaner single-channel signal. In this paper, the zeroth-order signal of the FOA signal is selected as the signal of the reference channel. The network of the neural beamformer is also based on SpatialNet with Mamba. Its structure is shown in Fig. 6. The mixed signal and the multi-channel separated signal are transformed to T-F domain through STFT, and then they are concatenated along the channel dimension, which is denoted as $\mathbf{Y_{cat}}\in\mathbb{R}^{F\times T\times 4M}$. The rest of the network structure of this neural beamformer is the same as that for target sound extraction, except that the number of input channels of the 1-D convolutional encoder for audio input is set to $4M$ ($2M$ for extracted multi-channel target signal, $2M$ for signal mixture), and the number of output channels of the 1-D deconvolutional layer is reduced to $2$ The neural beamformer still multiplies the framed intensity trajectory and the time-frequency domain signal after the encoders. The loss function of training the neural beamformer follows the one of target sound extraction network.
@@ IV-A1 Dataset generation
Our approach is validated using a simulated dataset sourced from FSD18k [43]. The dataset comprises 42 hours of signals sampled at 16kHz for training, 4 hours for validation and 4 hours for testing. Room dimensions are randomly selected from a range of $(3\times 3\times 3)m^{3}$ to $(10\times 10\times 6)m^{3}$, with reverberation times $T60$ randomly assigned between $0.2$s and $1.0$s. The microphone array is configured as an Eigenmike microphone array with 32 channels [44]. The captured signals are then converted into FOA format.
@@
The method for generating trajectories of the sound sources is same as that outlined in [45]. Initially, two points are randomly selected within the room: a starting point $\mathbf{p_{0}}=(p_{x0},p_{y0},p_{z0})$ and an endpoint $\mathbf{p_{N}}=(p_{xN},p_{yN},p_{zN})$. Connecting these two points establishes the fundamental trajectory line. Subsequently, sinusoidal trajectories with random frequencies $\mathbf{\omega}=(\omega_{x},\omega_{y},\omega_{z})$ and amplitudes $\mathbf{A}=(A_{x},A_{y},A_{z})$ are added along each axis. The number of oscillation periods is constrained to not exceed $2$. The position of the sound source at temporal sample $n$ can be calculated as:
$\begin{split}\mathbf{p_{n}}=\mathbf{p_{0}}+\frac{n}{N}(\mathbf{p_{N}}-\mathbf{p_{0}})+\mathbf{A}\cdot\sin(\mathbf{\omega}n),\end{split}$ (8)
@@
where $N$ is the length of the trajectory. In this experiment, only the DOA of the sound sources in utilized for target sound extraction. We set the number of sound sources to $2$. The RIRs is generated based on the Image Source Method [46] with the gpuRIR Python library [47]. Besides, additive Gaussian white sensor noise is added to audio mixtures with a uniformly sampled in the range of $[20,30]$ dB.
IV-A2 Network configuration
@@
The SpatialNet utilized consists of $8$ blocks [23]. The hidden unit size in each block is $96$. The kernel size of the 1-D convolutional encoders is $5$. The kernel size of the frequency-dimension group convolution is $3$. The group numbers are $8$. In Mamba, the dmodel parameter is set to $32$, state is set to $16$, dconv is set to $4$, and expand is set to $2$. The kernel size and stride for the max-pooling layer for calculating the envelope is $256$ samples (16 ms) and $128$ samples (8 ms) respectively. The STFT window length is $256$ samples (16 ms), the hop size is $128$ samples (8 ms), and the number of frequency bins is $129$. Each network in this paper is trained with Adam optimizer [48] with learning rate of $5\cdot 10^{-4}$.
IV-A3 Baselines
@@
we compare our method with Fasnet-TAC [12], which is a neural beamformer, and Inter-channel Conv-Tasnet (IC Conv-Tasnet) [17]. The end-to-end networks of SpatialNet with the same architecture of the network in this paper is also utilized as baseline. We also compare the results obtained using the SpatialNet architecture to train a multi-channel mapping network (MC-SpatialNet), which outputs the multi-channel signal of each source. From the multi-channel signals generated by MC-SpatialNet, we select the zero-th channel for comparison with final output of the proposed method. The code of proposed and baseline methods is available online [^1] https://github.com/WuDH2000/MFTS.
IV-A4 Metrics
The metrics utilized for evaluating the sound separation performance is SNR, signal to distortion ratio (SDR) and scale-invariant-signal-to-noise-ratio (SI-SNR). The performance of sound source tracking depends on the energy of the sound source; when the energy is low, the localization accuracy tends to deteriorate. On the other hand, the task of sound source separation focuses more on separating the source when it is active. Therefore, good localization results should aim for greater accuracy when the sound source is active, i.e., when the energy of the source is high. Based on this, we propose an Energy-Weighted Root Mean Square Angular Error (EWRMSAE) as an enhancement of the traditional RMSAE. The EWRMSAE () is defined as follows:
$\begin{split}&\text{EWRMSAE}=\\ &\sqrt{\frac{\sum_{n=0}^{N}(E^{trg}[n])^{2}\cdot(\text{AE}(\mathbf{traj}^{est}[n],\mathbf{traj}^{trg}[n]))^{2}}{\sum_{n=0}^{N}(E^{trg}[n])^{2}}},\end{split}$ (9)
where $AE(\cdot)$ calculates the absolute value of the angular error between the estimated and target trajectory at each temporal sample $n$. EWRMSAE assigns more weights to angular errors when sources have higher energy, while the weights of the angular errors when sources have lower energy are less.
@@ IV-A4 Metrics
Method SNR SI-SNR SDR
FasNet-TAC 10.81 10.03 11.71
IC Conv-Tasnet 11.83 10.92 12.60
SpatialNet 13.83 13.03 14.47
MC-SpatialNet 12.51 11.64 13.17
MFTS 15.35 14.48 15.83
Figure 7: SNR of different methods under different T60s.
Figure 8: SNR of deep learning-based methods under different angular separation between sources.
Figure 9: SNR at every 24 ms interval of deep learning-based methods under different angular separation between sources.
@@ IV-B Results
Table I shows each method’s average performance for source separation on the test dataset. The method proposed in this paper achieves the best performance. Compared to the baseline methods, our approach explicitly tracks the sound sources and utilizes the tracking results to aid separation. The improved separation results further enhance the precision of tracking, enabling the use of refined tracking results to facilitate better separation.
@@
To demonstrate the robustness of our proposed method, we compute the SNR values of all methods at different T60 values. Fig. 7 shows the results. It can be seen that our method maintains good performance under different reverberation times, and outperforms the baseline methods in all cases.
To investigate the impact of the spatial separation between sources on sound separation performance, we concurrently evaluate the method performance under stationary source conditions and conduct a statistical analysis of separation performance across varying angular separation between two sources. The experimental configuration mirrors the moving source scenario except for fixed source positions. Fig. 8 illustrates the SNR value versus angular separation between sources. The $0^{\circ}$ angular separation between sound sources indicates that they are simultaneously active at different distances in the same direction. It can be seen that the proposed method remains functional in stationary sound source scenarios. The proposed method demonstrates superiority over all baseline approaches across various source angular separations, except when compared to SpatialNet. When the angular separation between two sources exceeds $30^{\circ}$, the proposed approach begins to progressively outperform SpatialNet, with the performance gap widening as the angular separation increases, particularly beyond $90^{\circ}$, demonstrating effective spatial information exploitation. The performance of the proposed method and SpatialNet are almost the same when the angular separation between sound sources is below $30^{\circ}$, which can be attributed to the localization errors of the proposed method. When the angular separation between two sound sources falls below $30^{\circ}$ and the sources remain stationary, the localization errors inherent in the proposed method prevent effective differentiation between the two sources. However, this limitation becomes resolvable when sound sources exhibit movement. Even when the angular separation between sound sources falls below the localization precision threshold imposed by tracking errors, our framework compensates by leveraging temporal context from intervals with sufficient spatial separation. To validate this, we quantified the SNR value and the average angular separation over every 24-ms intervals under moving-source conditions, correlating separation performance with instantaneous angular separation between sources. Fig. 9 reveals that our method surpasses all baselines in short-term separation performance under moving source scenarios regardless of instantaneous angular separation, demonstrating the critical role of motion-induced temporal dynamics in resolving spatially ambiguous cases through temporal continuity constraints.
TABLE II: The NMSE of envelope estimation of the SpatialNet-based method and the NMF-based method.
Method NMSE
FastMNMF -3.36
SpatialNet -19.10
Figure 10: The EWRMSAE () under different T60s.
@@
Figure 11: The angular error () at individual audio samples under different amplitudes of the envelope signals.
IV-C Ablation study
To demonstrate the effectiveness of each part of the proposed method, we perform the ablation study. The ablation study is conducted under scenarios of two moving sources.
IV-C1 Initial tracking
We first check the effectiveness of the envelope estimation network. The proposed method based on SpatialNet is compared with the traditional signal processing method which uses FastMNMF for source separation followed by the envelope detector [49]. Table II shows the result. Table II demonstrates significant superiority of the SpatialNet-based method over traditional signal processing approaches.
To illustrate the efficiency of sound source tracking based on envelope estimation, we compare the performance of initial tracking network (init-track) to the method directly tracking the sources from the audio mixture (mix-track). Both of these two approaches utilize the same network, except that mix-track only takes the audio mixture as input, and outputs all the sources’ trajectories without utilizing the envelope. The loss function for training mix-track consists of MSE between the estimated trajectories and the differential loss. The uPIT [50] is used during training mix-track, which means that we perform the permutation on the entire trajectory and find the order of permutation with minimum value of the loss. Fig. 10 shows the EWRMSAE for both init-track and mix-track under different T60s. It can be observed that tracking based on envelope estimation results in a lower EWRMSAE across different T60 values compared to tracking without envelope information.